OpenStudy (anonymous):

Solve the following system of equations. x + y = 14 -x + y = 16

jimthompson5910 (jim_thompson5910):

Add the equations to get x + y = 14 -x + y = 16 ----------- 0x + 2y = 30 so 2y = 30 y = ???

jimthompson5910 (jim_thompson5910):

when you find y, use it to find x